OpenStudy (anonymous):

suppose you draw a card form a well- shuffled deck of cards 52 cards.What is the probablity of drawing a 7 or a jack? p(7 or a jack)= p(7)+ p( jack)=?

OpenStudy (m):

8/52

OpenStudy (anonymous):

4 7s and 4 jacks = 8 possibilities out of 52 cards = 8/52 = 2/13